# Latent Reasoning in LLMs as a Vocabulary-Space Superposition

### v1

@@ 6 Analysis: Understanding the Essence of Latent Reasoning
To quantify how much explicit information is preserved at each step of latent reasoning, we introduce the Top-K Effective Compression Ratio (ECR@K). This metric measures the proportion of explicit chain tokens that fall within the Top-K of the probability distribution at the corresponding latent step. Larger values indicate stronger horizontal compression across the reasoning chain, while values greater than one reflect true compression of the explicit chain. The formal definition is provided in the Appendix D.2.
@@
Figure 6: $N_{\mathrm{eff}}$ values of Latent-SFT on the Multi-Chain GSM8k-Aug dataset. The legend reports the mean values of both $N_{\mathrm{eff}}$ and the Top-2 Score.
Hypothesis 2. Latent Reasoning Represents the Superposition of Multiple Reasoning Paths.
To examine whether latent reasoning corresponds to the superposition of multiple paths, we define the effective global parallelism (Neff), which quantifies global evidence for the presence of multiple parallel reasoning paths. A larger value reflects a higher degree of parallelism. In addition, we introduce the Top-2 Score, defined as the ratio of the probabilities of the two most likely paths. Formal definitions are provided in the Appendix D.3. To assess the parallel reasoning capability of latent models, we construct multiple alternative reasoning chains corresponding to a single explicit reasoning path within GSM8k-Aug dataset (Multi-Chain GSM8k-Aug dataset in Appendix C).
@@ 7 Conclusion
In this work, we propose Latent-SFT, a method that enables latent reasoning in LLMs by defining latent tokens within the column space of the vocabulary. Experimental results show that on low-difficulty mathematical reasoning datasets, Latent-SFT achieves performance comparable to explicit COT-SFT while significantly shortening reasoning chains. On high-difficulty datasets, it outperforms methods based on last hidden state representations. Furthermore, our analysis demonstrates that latent reasoning embodies both the compression of a single reasoning chain and the superposition of multiple chains.
8 Ethics statement and Reprodicibility statement
This work does not raise ethical concerns. The code, datasets, and models of our method are available in an GitHub repository https://github.com/DJC-GO-SOLO/Latent-SFT.
@@ Appendix A Use of LLMs
Appendix B Implementation Details
Since Latent-SFT requires weighting with respect to the vocabulary matrix, we freeze the vocabulary parameters during training. All models are trained using LoRA with rank 64. Following COCONUT, we initialize all methods from the COT-SFT model to accelerate convergence. In practice, we find that a fixed-length segmentation function performs better than a semantic-based segmentation. Therefore, all results reported in this paper are based on fixed-length segmentation functions. For LLaMA-3.2-1B-Instruct, we set the learning rate to $1\times 10^{-4}$ in Step 1 and $4\times 10^{-4}$ in Step 2. For DeepSeek-Distill-Qwen-7B, the learning rates are $5\times 10^{-5}$ (Step 1) and $1\times 10^{-5}$ (Step 2). Following DeepSeek-AI & et al. (2025), we use a decoding temperature of 0.6 and top-p of 0.95. We did not perform parameter tuning for $\lambda$ and $\beta$; setting both to 1 yielded satisfactory performance.
During evaluation, low-difficulty mathematical reasoning datasets were inferred using the Transformers library with a maximum sequence length of 128. High-difficulty datasets were evaluated with the improved SGlang library, supporting a maximum token length of 32,768. All training and inference experiments were conducted on 8×A100 or 8×H100 GPUs.
Appendix C Construction of the Multi-Chain GSM8k-Aug Dataset
@@
Figure 8: An example of calling GPT5 and post-processing.
Appendix D Analytical Metrics
D.1 Notation and Preliminaries
@@
D.2 ECR@K (Effective Compression Rate)
@@ D.3 Neff and Top-2 Score
Figure 10: An example of Latent-SFT(2) reasoning on the GSM8k-Aug test set. The reasoning chain length is reduced by nearly half.

### v2

@@ 7 Essence of Latent Reasoning
To quantify how much explicit information is preserved at each step of latent reasoning, we introduce the Top-K Effective Compression Ratio (ECR@K). This metric measures the proportion of explicit chain tokens that fall within the Top-K of the probability distribution at the corresponding latent step. Larger values indicate stronger horizontal compression across the reasoning chain, while values greater than one reflect true compression of the explicit chain. The formal definition is provided in the Appendix E.2.
@@
Hypothesis 2. Latent Reasoning as the Superposition of Multiple Reasoning Chains.
To examine whether latent reasoning corresponds to the superposition of multiple paths, we define the effective global parallelism (Neff), which quantifies global evidence for the presence of multiple parallel reasoning paths. A larger value reflects a higher degree of parallelism. In addition, we introduce the Top-2 Score, defined as the ratio of the probabilities of the two most likely paths. Formal definitions are provided in the Appendix E.3. To assess the parallel reasoning capability of latent models, we construct multiple alternative reasoning chains corresponding to a single explicit reasoning path within GSM8k-Aug dataset (Multi-Chain GSM8k-Aug dataset in Appendix D).
@@ 8 Conclusion
In this work, we introduce Latent-SFT, a unified framework designed to enable latent reasoning to realize the ideal state of multi-path superposition. By systematically addressing challenges at the token, chain, and learning levels, we transform latent reasoning from an unconstrained ”black box” into a structured, interpretable process. Empirical evaluations across six mathematical benchmarks demonstrate that Latent-SFT consistently outperforms explicit SFT while achieving a 2.7x to 5.5x reduction in reasoning chain length. Crucially, our analysis confirms that the learned latent process is not merely a compression of a single explicit path, but a superposition of multiple reasoning chains. This work suggests that the future of efficient LLM reasoning lies not in discarding the chain-of-thought, but in evolving it into a multiple paths, high-entropy latent reasoning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Use of LLMs
Appendix B Training Computational Costs
We provide a detailed breakdown of the computational time required for both the Latent-Chain Construction and Latent-Optim phases. All models were trained on 8 $\times$ NVIDIA A100 GPUs.
The training process for Latent-SFT is divided into two primary steps. The first step, Latent-Chain Construction, involves three sub-phases: (1) Encoder Training, (2) Decoder Training, and (3) Joint Fine-tuning. The second step, Latent-Optim, involves training the LLM using the proposed stochastic Gumbel-Softmax strategy.
Table 3 summarizes the training duration for each model size. Despite the multi-stage nature of our framework, the total training cost remains manageable.
Table 3: Training time breakdown (in hours) for Latent-SFT across different model scales.
Base Model Latent-Chain Construction (h) Latent-Optim (h) Total (h)
Encoder Decoder Joint Subtotal
LLaMA-3.2-1B-Instruct 4.0 4.0 4.5 12.5 12.0 24.5
LLaMA-3.2-3B-Instruct 6.5 6.5 8.0 21.0 22.0 43.0
DeepSeek-Distill-Qwen-7B 12.0 12.0 15.0 39.0 50.0 89.0
Appendix C Theoretical Analysis of Stochastic Latent Optimization
In this section, we analyze the implicit regularization effect of Gumbel-noise injection. Unlike standard Gaussian noise, Gumbel noise has a non-zero mean and fixed variance. We show that, combined with the translation invariance of Softmax, this induces a Hessian-based regularization term in expectation.
Theorem 1 (Implicit Hessian Regularization via Gumbel-Softmax). Let $\bm{\alpha}\in\Delta^{V-1}$ be a static target distribution in the simplex interior (i.e., $\alpha_{i}>0$ and $\sum_{i}\alpha_{i}=1$). Consider the stochastic target generated via Eq. (6): $\tilde{\mathbf{p}}=\sigma\!\left((\log\bm{\alpha}+\mathbf{g})/\tau\right)$, where $\sigma(\cdot)$ is Softmax, $\tau>0$, and $\mathbf{g}\in\mathbb{R}^{V}$ has i.i.d. entries $g_{i}\sim\mathrm{Gumbel}(0,1)$. Define the unperturbed baseline
$\mathbf{p}_{\tau}\;\triangleq\;\sigma\!\left(\tau^{-1}\log\bm{\alpha}\right).$ (8)
We analyze the expected loss in a neighborhood of $\mathbf{p}_{\tau}$ under the small-perturbation regime (defined below). Then, for a twice-differentiable loss $\mathcal{L}(\theta,\mathbf{p})$ that depends on $\mathbf{p}$ through the forward computation (hence generally non-linear in $\mathbf{p}$), we have:
$\mathbb{E}_{\mathbf{g}}\!\left[\mathcal{L}(\theta,\tilde{\mathbf{p}})\right]=\mathcal{L}(\theta,\mathbf{p}_{\tau})+\frac{\pi^{2}}{12\tau^{2}}\,\mathrm{Tr}\!\Big(\mathbf{J}_{\sigma}(\mathbf{p}_{\tau})^{\top}\,\nabla_{\mathbf{p}}^{2}\mathcal{L}(\theta,\mathbf{p})\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\,\mathbf{J}_{\sigma}(\mathbf{p}_{\tau})\Big)+\mathcal{O}\!\Big(\|\nabla_{\mathbf{p}}\mathcal{L}\|\,\mathbb{E}\|\bm{\delta}\|^{2}\Big)+\mathcal{O}\!\Big(\mathbb{E}\|\bm{\delta}\|^{3}\Big),$ (9)
where $\mathbf{J}_{\sigma}(\mathbf{p}_{\tau})=\mathrm{diag}(\mathbf{p}_{\tau})-\mathbf{p}_{\tau}\mathbf{p}_{\tau}^{\top}$ is the Softmax Jacobian evaluated at $\mathbf{p}_{\tau}$, and $\bm{\delta}$ is the effective (zero-mean) perturbation in logit space defined in the proof. In particular, in our experiments with $\tau=1$, $\mathbf{p}_{\tau}=\sigma(\log\bm{\alpha})=\bm{\alpha}$, so the expansion is centered at the semantic target. This indicates that Gumbel noise induces a curvature penalty that favors flatter neighborhoods around the target distribution, with leading-order scaling $1/\tau^{2}$.
Proof.
1. Notation and mean removal via translation invariance. Let $\mathbf{l}=\log\bm{\alpha}$. The stochastic target is
$\tilde{\mathbf{p}}=\sigma\!\left(\frac{\mathbf{l}+\mathbf{g}}{\tau}\right).$ (10)
A standard Gumbel variable has mean $\mathbb{E}[g]=\gamma$ and variance $\mathrm{Var}(g)=\pi^{2}/6$. Decompose $\mathbf{g}=\bm{\epsilon}+\gamma\mathbf{1}$ where $\mathbb{E}[\bm{\epsilon}]=\mathbf{0}$ and $\mathbb{E}[\bm{\epsilon}\bm{\epsilon}^{\top}]=\frac{\pi^{2}}{6}\mathbf{I}$. Using Softmax translation invariance $\sigma(\mathbf{x}+c\mathbf{1})=\sigma(\mathbf{x})$, we obtain
$\tilde{\mathbf{p}}=\sigma\!\left(\frac{\mathbf{l}+\bm{\epsilon}}{\tau}\right)=\sigma\!\left(\frac{\mathbf{l}}{\tau}+\bm{\delta}\right),\qquad\bm{\delta}\triangleq\frac{\bm{\epsilon}}{\tau}.$ (11)
Define the baseline $\mathbf{p}_{\tau}\triangleq\sigma(\mathbf{l}/\tau)$, so that $\tilde{\mathbf{p}}=\sigma(\mathbf{z}_{\tau}+\bm{\delta})$ with $\mathbf{z}_{\tau}=\mathbf{l}/\tau$ and $\tilde{\mathbf{p}}(\mathbf{0})=\mathbf{p}_{\tau}$.
2. Second-order expansions. We assume a small-perturbation regime where $\|\bm{\delta}\|$ is sufficiently small such that second-order Taylor expansions are valid and the remainders are controlled by $\mathcal{O}(\|\bm{\delta}\|^{3})$ in expectation.
Softmax expansion. Let $\mathcal{H}_{\sigma}$ denote the (third-order) Hessian tensor of Softmax with respect to logits. Using tensor contraction notation $\mathcal{H}_{\sigma}[\bm{\delta},\bm{\delta}]\in\mathbb{R}^{V}$,
$\tilde{\mathbf{p}}-\mathbf{p}_{\tau}=\mathbf{J}_{\sigma}(\mathbf{p}_{\tau})\,\bm{\delta}+\frac{1}{2}\,\mathcal{H}_{\sigma}[\bm{\delta},\bm{\delta}]+\mathcal{O}(\|\bm{\delta}\|^{3}).$ (12)
Loss expansion. Expanding $\mathcal{L}(\theta,\mathbf{p})$ around $\mathbf{p}=\mathbf{p}_{\tau}$,
[…]
Substituting the first-order relation $\tilde{\mathbf{p}}-\mathbf{p}_{\tau}=\mathbf{J}_{\sigma}\bm{\delta}+\mathcal{O}(\|\bm{\delta}\|^{2})$ into the quadratic term yields
$\mathcal{L}(\theta,\tilde{\mathbf{p}})=\mathcal{L}(\theta,\mathbf{p}_{\tau})+\left(\mathbf{J}_{\sigma}\bm{\delta}+\frac{1}{2}\mathcal{H}_{\sigma}[\bm{\delta},\bm{\delta}]\right)^{\top}\nabla_{\mathbf{p}}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}+\frac{1}{2}\bm{\delta}^{\top}\mathbf{J}_{\sigma}^{\top}\nabla_{\mathbf{p}}^{2}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\mathbf{J}_{\sigma}\,\bm{\delta}+\mathcal{O}(\|\bm{\delta}\|^{3}),$ (14)
where $\mathbf{J}_{\sigma}=\mathbf{J}_{\sigma}(\mathbf{p}_{\tau})$ for brevity.
3. Taking expectations. Since $\mathbb{E}[\bm{\delta}]=\mathbf{0}$, the linear-in-$\bm{\delta}$ term vanishes: $\mathbb{E}[\mathbf{J}_{\sigma}\bm{\delta}]=\mathbf{0}$. The remaining first-order bias arises from Softmax nonlinearity:
$\mathbb{E}\!\left[\left(\tfrac{1}{2}\mathcal{H}_{\sigma}[\bm{\delta},\bm{\delta}]\right)^{\top}\nabla_{\mathbf{p}}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\right]=\mathcal{O}\!\Big(\|\nabla_{\mathbf{p}}\mathcal{L}\|\,\mathbb{E}\|\bm{\delta}\|^{2}\Big),$ (15)
which becomes negligible near a stationary point where $\|\nabla_{\mathbf{p}}\mathcal{L}\|\approx 0$.
For the quadratic term,
$\mathbb{E}\!\left[\frac{1}{2}\bm{\delta}^{\top}\mathbf{J}_{\sigma}^{\top}\nabla_{\mathbf{p}}^{2}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\mathbf{J}_{\sigma}\,\bm{\delta}\right]$ $=\frac{1}{2}\,\mathrm{Tr}\!\left(\mathbf{J}_{\sigma}^{\top}\nabla_{\mathbf{p}}^{2}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\mathbf{J}_{\sigma}\,\mathbb{E}[\bm{\delta}\bm{\delta}^{\top}]\right).$ (16)
Using $\mathbb{E}[\bm{\epsilon}\bm{\epsilon}^{\top}]=\frac{\pi^{2}}{6}\mathbf{I}$ and $\bm{\delta}=\bm{\epsilon}/\tau$ gives $\mathbb{E}[\bm{\delta}\bm{\delta}^{\top}]=\frac{\pi^{2}}{6\tau^{2}}\mathbf{I}$, hence
$\mathbb{E}\!\left[\frac{1}{2}\bm{\delta}^{\top}\mathbf{J}_{\sigma}^{\top}\nabla_{\mathbf{p}}^{2}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\mathbf{J}_{\sigma}\,\bm{\delta}\right]=\frac{\pi^{2}}{12\tau^{2}}\,\mathrm{Tr}\!\Big(\mathbf{J}_{\sigma}^{\top}\nabla_{\mathbf{p}}^{2}\mathcal{L}\big|_{\mathbf{p}=\mathbf{p}_{\tau}}\mathbf{J}_{\sigma}\Big).$ (17)
Collecting terms yields Eq. (9), with the remainder $\mathcal{O}(\mathbb{E}\|\bm{\delta}\|^{3})$ from Taylor truncation. $\square$
Appendix D Construction of the Multi-Chain GSM8k-Aug Dataset
@@
Figure 8: Visualization of last-layer hidden states and token embeddings of the Deepseek-Distill-Qwen-7B model on the Math500 dataset.
Figure 9: Illustration of the Multi-Chain dataset construction pipeline. We prompt GPT-5 to generate alternative valid reasoning paths for a given GSM8k problem. The response section demonstrates the diversity filtering mechanism: candidate paths are evaluated based on edit distance, and those exhibiting high similarity to the standard solution or previously accepted paths (marked with $\times$) are rejected to ensure significant reasoning diversity.
Appendix E Analytical Metrics
E.1 Notation and Preliminaries
@@
E.2 ECR@K (Effective Compression Rate)
@@ E.3 Neff and Top-2 Score
Figure 10: Reasoning example of Latent-SFT(2) on the GSM8k-Aug test set. The model generates a sequence of latent tokens (annotated as ”Top-1 Tokens”) instead of explicit natural language steps. While these tokens appear as unintelligible numeric combinations to humans, they effectively encode the necessary logical computations, allowing the model to derive the correct answer 18. Notably, this latent process reduces the reasoning chain length by nearly 50% compared to the reference solution, demonstrating high computational efficiency.